OpenStudy (anonymous):

Find all positive integer pairs (x,y) such that 4xy-x-y is a perfect square.

OpenStudy (anonymous):

without loss of generality we can suppose \[y\ge x \]because the expression is symmetric

OpenStudy (anonymous):

assume that \[y=x+a\]with \(a\ge 0\)

OpenStudy (anonymous):

Okay! I'll try from here!!

OpenStudy (anonymous):

wait ... im afraid i made a mistake :/

OpenStudy (anonymous):

Don't be afraid :P \[4xy-x-y=4x(x+a)-x-(x+a)=4x^2+(4a-2)x-a\]

OpenStudy (anonymous):

\[\Delta=(4a-2)^2 - 4(4)(-a)=0\]\[16a^2 - 16a + 4 + 16a=0\]\[16a^2 + 4 =0\]Well well well...

OpenStudy (shubhamsrg):

4xy-x-y = (x+y)^2 - (x-y)^2 - (x+y) = (x+y)(x+y-1) - (x-y)^2 i am not too sure if that'd help! :P

OpenStudy (anonymous):

@RolyPoly ...Are you familiar with modular arithmetic??

OpenStudy (anonymous):

I've never heard of it! :(

OpenStudy (shubhamsrg):

i dont know much but i do have a little idea..will you please guide me/us through a solution @mukushla

OpenStudy (anonymous):

Forgive me for closing this question, even though it's not yet solved!